40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 15749     Entry time: Wed Jan 6 16:18:38 2021
Author: gautam 
Type: Update 
Category: Optical Levers 
Subject: BS Oplev glitchy 

As part of the hunt why the X arm IR transmission RIN is anomalously high, I noticed that the BS Oplev Servo periodically kicks the optic around - the summary pages are the best illustration of this happening. Looking back in time, these seem to have started ~Nov 23 2020. The HeNe power output has been degrading, see Attachment #1, but this is not yet at the point where the head usually needs replacing. The RIN spectrum doesn't look anomalous to me, see Attachment #2 (the whitening situation for the quadrants is different for the BS and the TMs, which explains the HF noise). I also measured the loop UGFs (using swept sine) - seems funky, I can't get the same coherence now (live traces) between 10-30 Hz that I could before (reference traces) with the same drive amplitude, and the TF that I do measure has a weird flattening out at higher frequencies that I can't explain, see Attachment #3.

The excess RIN is almost exactly in the band that we expect our Oplevs to stabilize the angular motion of the optics in, so maybe needs more investigation - I will upload the loop suppression of the error point later. So far, I don't see any clean evidence of the BS Oplev HeNe being the culprit, so I'm a bit hesitant to just swap out the head...

Attachment 1: missingData.png  368 kB  | Hide | Hide all
missingData.png
Attachment 2: OLRIN.pdf  70 kB  | Hide | Hide all
OLRIN.pdf
Attachment 3: BS_OL_P.pdf  16 kB  | Hide | Hide all
BS_OL_P.pdf
Attachment 4: BS_OL_suppression.pdf  108 kB  Uploaded Wed Jan 6 18:28:26 2021  | Hide | Hide all
BS_OL_suppression.pdf
ELOG V3.1.3-